OpenStudy (anonymous):

Angle U and Angle V are complementry angles. If Angle U is three less than twice Angle V What are the measures of the Angles

OpenStudy (ganpat):

U = (180 - V).... as they are complementary.. U = 2V - 3 180 - V = 2V - 3 177 = 3V V = 59... SO U = 180 - 59 = 121

OpenStudy (anonymous):

Thank you

OpenStudy (ganpat):

w/c..

OpenStudy (anonymous):

is that the same for V

OpenStudy (ganpat):

what V ?? if u inter change U and V .. u will find interchanged answers..

OpenStudy (anonymous):

thought so thank you just double checking

OpenStudy (ganpat):

anytime..